# The Gaussian-Multinoulli Restricted Boltzmann Machine A Potts Model Extension of the GRBM

### v1

@@ Abstract
Many real-world tasks, from associative memory to symbolic reasoning, demand discrete, structured representations that standard continuous latent models struggle to express naturally. We introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM), a generative energy-based model that extends the Gaussian-Bernoulli RBM (GB-RBM) by replacing binary hidden units with $q$-state Potts variables. This modification enables a combinatorially richer latent space and supports learning over multivalued, interpretable latent concepts. We formally derive GM-RBM’s energy function, learning dynamics, and conditional distributions, showing that it preserves tractable inference and training through contrastive divergence. Empirically, we demonstrate that GM-RBMs model complex multimodal distributions more effectively than binary RBMs, outperforming them on tasks involving analogical recall and structured memory. Our results highlight GM-RBMs as a scalable framework for discrete latent inference with enhanced expressiveness and interoperability.
@@ 1 Introduction
Restricted Boltzmann Machines (RBMs) constitute a class of undirected energy-based models defined on a bipartite graph between visible units $\mathbf{v}$ and latent units $\mathbf{h}$(hidden). Their tractable block Gibbs updates [6] and scalable approximate inference via contrastive divergence enable efficient training on high-dimensional datasets [1, 6]. In the binary hidden configuration, an RBM reduces to an Ising model where hidden units $h_{j}\in{-1,+1}$ interact pairwise through couplings $W_{ij}$, yielding an energy landscape analogous to classical spin systems [14]. The Gaussian–Bernoulli RBM (GB-RBM) augments this formulation by replacing these visible variables with Gaussian potentials, while preserving binary hidden spins [7]. Although GB-RBMs capture continuous inputs such as image intensities or sensor readings, the binary latent constraints limit their expressivity when representing multimodal or discrete feature correlations [4, 15] .
The $q$-state Potts model generalizes the Ising model by allowing each spin $s_{j}$ to occupy one of $q$ discrete states, enriching the model’s capacity to encode complex interaction patterns and phase transitions [22, 32]. Potts-type energy functions have demonstrated efficacy in image segmentation, combinatorial optimization, community detection, and categorical embedding tasks, where capturing higher-order discrete structures is key [12]. Building on these principles, we introduce the Gaussian–Multinoulli Restricted Boltzmann Machine (GM-RBM), which integrates Gaussian-distributed visible units with a hidden layer of $q$-state Potts spins. This GB-RBM-like architecture provides a clear path to benchmarking and showcasing the power of Potts-based energy functions.
The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) we propose has a small change in the underlying representation of hidden nodes that has a profound improvement on model performance. While prior work has addressed mixed-variate inputs using Gaussian-Bernoulli structures [26], convolutional and tiled weight-sharing schemes for spatial hierarchies [10], and implicit mixture models of RBMs for multimodal density estimation [20], these approaches retain binary or sigmoidal-based latent representations, limiting their ability to capture categorical variation at the feature level. In contrast, the GM-RBM replaces binary hidden units with $q$-state Potts (Multinoulli) variables, enabling each hidden unit to express one of several discrete latent modes rather than mere presence or absence. This yields a latent space that scales as $q^{m}$ (with $m$ hidden units), and facilitates more structured, interpretable, and compositional generative behavior. Unlike energy-based models that treat discrete latents as auxiliary sampling artifacts [4], or continuous latent embeddings for mixed data [25], the GM-RBM directly leverages categorical latent dynamics to build rich distributions over continuous-valued visible units. This hybrid formulation opens new avenues for modeling structured analog data, such as trajectories, natural images, or biophysical time series, with discrete control over generative features.
2 Background and Model Overview
2.1 Background and Motivation
Restricted Boltzmann Machines (RBMs) have played a foundational role in energy-based modeling and unsupervised representation learning. In particular, Gaussian-Bernoulli RBMs (GB-RBMs), which pair real-valued Gaussian visible units with binary latent variables, have been adopted for modeling continuous input data such as images and signals [15]. However, while GB-RBMs are capable of learning low-level features, they face inherent limitations in their ability to capture structured, discrete latent representations.
The root of this limitation lies in the mismatch between the binary latent space and the Gaussian visible manifold. The Gaussian prior enforces a continuous energy landscape in the input space, which, when combined with binary latent units, leads to the learning of clustered, overlapping encodings. These encodings tend to coalesce into fuzzy representations that lack the expressiveness needed for capturing discrete abstractions, such as object parts, symbolic states, or multimodal feature variants.
Moreover, the binary latent units in GB-RBMs act as on/off switches, providing limited control over the diversity and structure of the hidden space. This bottleneck hampers the model’s ability to disentangle categorical or compositional structure, particularly when the data exhibits rich semantic variation.
2.2 Motivating Discrete Latent Structure: Potts Units
In contrast, Potts models, which extend binary variables to q-state categorical variables, offer a compelling mechanism for reintroducing structured discrete latent representations into energy-based models. By allowing each hidden unit to express one of multiple mutually exclusive states, Potts units can represent discrete modes, templates, or symbolic features directly. This increases the combinatorial capacity of the model from $2^{m}$ (binary) to $q^{m}$ (categorical), enabling richer and more interpretable feature representations.
Importantly, Potts units also induce a sharper factorization of latent structure, allowing for the emergence of compositional encodings where different latent units capture orthogonal or complementary concepts. This categorical structure aligns more naturally with real-world semantics and improves both generalization and controllability.
2.3 The Gaussian-Multinoulli RBM (GM-RBM)
To address the limitations of GB-RBMs and exploit the representational advantages of Potts variables, we introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM). In this architecture, the visible layer consists of Gaussian-valued units, modeling real-valued data such as sensor signals or pixel intensities while the hidden layer consists of Multinoulli (Potts) units, each of which selects one of q discrete latent states.
Each Potts state corresponds to a different “template” or filter that contributes to the energy of the model, allowing a single hidden unit to encode latent modes or feature classes rather than just presence/absence. The GM-RBM thus combines the smooth representational power of Gaussian inputs with the symbolic richness of discrete latent categories.
This hybrid construction restores the model’s ability to capture structured, discrete latent features while retaining flexibility in modeling continuous-valued observations. As we show, the GM-RBM can learn more interpretable latent spaces, support combinatorial recombination of features, and exhibit sharper mode separation in the visible manifold than its Gaussian-Bernoulli counterpart.
3 Theoretical Foundations of the Gaussian-Multinoulli RBM (GM-RBM)
The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) extends the classical Gaussian-Bernoulli RBM by replacing binary latent units with discrete $q$-state Potts (Multinoulli) variables. This modification introduces a combinatorially structured latent space capable of expressing discrete modes while modeling continuous-valued data in the visible layer. In this section, we develop a mathematical understanding of how this architecture supports richer and more structured representations.
3.1 Energy Function and Architecture
The GM-RBM defines a joint energy over visible variables $v\in\mathbb{R}^{n}$ and hidden Potts variables $h=(h_{1},\dots,h_{m})$, where each $h_{j}\in\{1,\dots,q\}$ selects a discrete state for hidden unit $j$:
$E(v,h)=\frac{1}{2}\sum_{i=1}^{n}(v_{i}-b_{i})^{2}-\sum_{j=1}^{m}c_{j,h_{j}}-\sum_{i=1}^{n}\sum_{j=1}^{m}W_{ij}^{(h_{j})}v_{i}$ (1)
This can be re-expressed in terms of the conditional Gaussian mean $\mu(h)$ as:
$E(v,h)=\frac{1}{2}\|v-\mu(h)\|^{2}+\text{const}(h),\quad\text{where}\quad\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (2)
Each hidden unit contributes a state-specific template vector $W_{j}^{(h_{j})}$, and the visible layer is modeled as a Gaussian centered at the compositional mean $\mu(h)$:
$P(v\mid h)=\mathcal{N}(\mu(h),I)$ (3)
3.2 Compositionality and Factored Representation
The mean vector $\mu(h)$ is a sum over learned templates selected by the Potts states:
$\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (4)
This creates a structured and factored representation in visible space. Each hidden unit encodes one of $q$ possible modes, and the overall visible configuration is a compositional sum of these latent components. This modularity enables reuse of latent features across samples, disentangled latent representation, and combinatorial expressivity across $q^{m}$ configurations
The set of all compositional means $\mu(h)$ forms a structured lattice in $\mathbb{R}^{n}$. The resulting energy landscape has the following properties:
  • Locally linear: Each mode defines a Gaussian with fixed covariance
  • Globally discrete: The mode structure is defined by a combinatorial Potts lattice
  • Modular: Latent units contribute independently to the visible pattern
This structure allows GM-RBMs to model multimodal distributions over continuous space with precise symbolic control via latent discrete selections.

### v2

# The Gaussian-Multinoulli Restricted Boltzmann Machine: A Potts Model Extension of the GRBM

@@ Abstract
Many real-world tasks, from associative memory to symbolic reasoning, benefit from discrete, structured representations that standard continuous latent models can struggle to express. We introduce the Gaussian–Multinoulli Restricted Boltzmann Machine (GM-RBM), a generative energy-based model that extends the Gaussian–Bernoulli RBM (GB-RBM) by replacing binary hidden units with $q$-state categorical (Potts) units, yielding a richer latent state space for multivalued concepts. We provide a self-contained derivation of the energy, conditional distributions, and learning rules, and detail practical training choices (contrastive divergence with temperature annealing and intra-slot diversity constraints) that avoid state collapse. To separate architectural effects from sheer latent capacity, we evaluate under both capacity-matched and parameter-matched setups, comparing GM-RBM with GB-RBM configured to have the same number of possible latent assignments. On analogical recall and structured memory benchmarks, GM-RBM achieves competitive, and in several regimes, improved recall at equal capacity with comparable training cost, despite using only Gibbs updates. The discrete $q$-ary formulation is also amenable to efficient implementation. These results clarify when categorical hidden units provide a simple, scalable alternative to binary latents for discrete inference within tractable RBMs.
@@ 1 Introduction
Restricted Boltzmann Machines (RBMs) are energy-based models with an undirected bipartite graph structure with visible and hidden units and a subset of the broader fully connected Boltzmann machines. The lack of intralayer connections enables training tractability through parallel updates of all units within each layer (Block Gibbs updates) through a biased stochastic gradient estimator for log-likelihood (Contrastive Divergence) (Hinton, 2002; Hinton and Salakhutdinov, 2006). Although powerful, the strictly binary units in hidden and visible units make them non-ideal for dealing with multivalued data. One approach to extend their use to continuous data is to replace the visible units with Gaussian units, termed Gaussian–Bernoulli RBMs (GB-RBMs) (Liao et al., 2022). However, their binary hidden units often struggle with inherently categorical, mutually exclusive factors. We address this mismatch by replacing each binary hidden unit with a one-of-$q$ categorical unit (Potts), yielding the Gaussian–Multinoulli RBM (GM-RBM): Gaussian visibles paired with $q$-state latent hidden units. Conceptually, this keeps the RBM’s simplicity but aligns inductive bias with categorical structure, effectively allowing the model to capture inherent, underlying categorical relationships. Prior work tended to retain binary latents as they fit existing tooling and samplers; in contrast, categorical slots raise practical issues (e.g., intra-slot degeneracy) and make fair comparisons challenging. In order to separate architectural effects from raw capacity, we design comparison protocols and clarify the challenges associated with them.
Our contributions are threefold: (i) a drop-in Potts hidden layer that preserves tractable conditionals while retaining the standard RBM training pipeline; (ii) comparison protocols (capacity-matched and parameter-matched) that isolate the effect of categorical slots; and (iii) empirical results showing that, with equal negative-phase budgets and pure block Gibbs updates, increasing $q$ consistently improves image quality (FID) and hetero-associative recall, closing much of the gap to GB-RBMs that require costlier Gibbs–Langevin steps (Liao et al., 2022). The key message being a minimal architectural change yields disproportionate gains due to the q-slot augmented latent: sharper posteriors, more interpretable codes, and stronger retrieval and generation while utilizing similar compute resources.
2 Background and Motivation
2.1 Background
Boltzmann machines (BM) are a class of energy-based models utilizing binary units ($\{-1,1\}$) that sample from a Boltzmann probability distribution over $x=(v,h)$ given by:
$p_{\theta}(x)=\frac{1}{Z(\theta)}\exp\!\big(-E_{\theta}(x)\big),\qquad Z(\theta)=\sum_{x}\exp\!\big(-E_{\theta}(x)\big).$
where $\theta$ is a set of biases ($a_{i}$) and weights ($W_{ij}$). (Smolensky, 1986; Hinton, 2002) The energy function for a BM is given by:
$E_{\theta}(x)=-\sum_{i}a_{i}x_{i}-\tfrac{1}{2}\sum_{i\neq j}W_{ij}x_{i}x_{j},\qquad W_{ij}=W_{ji},\;W_{ii}=0.$
Log-likelihood learning decomposes into a positive (data) and negative (model) phase:
$\frac{\partial}{\partial\theta}\log p_{\theta}(v)=\mathbb{E}_{p(h\mid v)}\!\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big]-\mathbb{E}_{p(v\mid h)}\!\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big],$
which yields the classic correlation-matching rules:
$\frac{\partial}{\partial W_{ij}}\log p_{\theta}(v)=\langle x_{i}x_{j}\rangle_{\text{data}}-\langle x_{i}x_{j}\rangle_{\text{model}},\qquad\frac{\partial}{\partial a_{i}}\log p_{\theta}(v)=\langle x_{i}\rangle_{\text{data}}-\langle x_{i}\rangle_{\text{model}}.$
However, correlation-matching rules are intractable for most real-world data (Liao et al., 2022). One way to overcome this is to use the Restricted Boltzmann machine: Restricted Boltzmann machines split the graph into a bipartite structure with visible units $v$ and hidden units $h$, removing within-layer edges so that the joint probability distribution can be broken down into simpler conditional probabilities enabling block Gibbs sampling. For binary visibles $v\in\{0,1\}^{n}$ and binary hiddens $h\in\{0,1\}^{m}$,
$E(v,h)=-b^{\top}v-c^{\top}h-v^{\top}Wh,$
$p(h_{j}=\textbf{1}\mid v)=\sigma\big((W^{\top}v)_{j}+c_{j}\big),\qquad p(v_{i}=\textbf{1}\mid h)=\sigma\big((Wh)_{i}+b_{i}\big),$
and the learning updates specialize to moment matching between the two layers,
$\Delta W\propto\mathbb{E}[vh^{\top}]_{\text{data}}-\mathbb{E}[vh^{\top}]_{\text{model}},\quad\Delta b\propto\mathbb{E}[v]_{\text{data}}-\mathbb{E}[v]_{\text{model}},\quad\Delta c\propto\mathbb{E}[h]_{\text{data}}-\mathbb{E}[h]_{\text{model}}.$
Contrastive divergence(CD) approximates model expectations by alternating $h\!\sim\!p(h\mid v)$ and $v\!\sim\!p(v\mid h)$ for a $k$ steps of data CD-$k$ or for a persistence chain (persistent CD). Usually, both hidden and visible units are taken as binary, but for continuous data it is natural to use Gaussian visible units (Liao et al., 2022). In the Gaussian–Bernoulli RBM (Hinton), visibles are Gaussian with diagonal variance $\sigma^{2}$ and hiddens remain binary. A convenient parameterization is:
$E(v,h)=\sum_{i}\frac{(v_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}-\sum_{i,j}\frac{v_{i}}{\sigma_{i}^{2}}W_{ij}h_{j}-\sum_{j}b_{j}h_{j},$
which gives closed-form conditionals used inside CD,
$p(v\mid h)=\mathcal{N}\!\big(\mu+Wh,\ \mathrm{diag}(\sigma^{2})\big),\qquad p(h_{j}=1\mid v)=\text{Sigmoid}\!\Big(\big[W^{\top}(v\oslash\sigma^{2})\big]_{j}+b_{j}\Big).$
The corresponding stochastic-gradient updates keep the same positive/negative structure but respect the visible scaling,
$\Delta\bm{W}\propto\Big\langle\left(\frac{\mathbf{v}}{\bm{\sigma}^{2}}\right)\,\mathbf{h}^{\top}\Big\rangle_{\text{data}}-\Big\langle\left(\frac{\mathbf{v}}{\bm{\sigma}^{2}}\right)\,\mathbf{h}^{\top}\Big\rangle_{\text{model}},$
$\Delta\bm{\mu}\propto\Big\langle\frac{\mathbf{v}-\bm{\mu}}{\bm{\sigma}^{2}}\Big\rangle_{\text{data}}-\Big\langle\frac{\mathbf{v}-\bm{\mu}}{\bm{\sigma}^{2}}\Big\rangle_{\text{model}}$
$\Delta\bm{b}\propto\langle\bm{h}\rangle_{\text{data}}-\langle\bm{h}\rangle_{\text{model}}\qquad$
In this work we extend the role of the binary hidden variables by replacing each Bernoulli with a one-of-$q$ categorical (Potts) slot, aligning the latent prior with mutually exclusive factors while preserving the RBM’s block-sampling tractability and the Gaussian visible layer.
2.2 Motivating Principles
Many perceptual and symbolic factors are naturally categorical and mutually exclusive. Approximating such structure with many Bernoulli latents (as in a GB–RBM) forces variants to be represented by co–activating subsets of units, which encodes information across the hidden layer and yields ambiguous codes.
The Gaussian–Multinoulli RBM (GM–RBM) encodes each factor as a one-of-$q$ slot. With $m$ slots and hidden configuration $h=(h_{1},\dots,h_{m})$ where $h_{j}\in\{1,\dots,q\}$, the visible conditional remains Gaussian,
$p(v\mid h)=\mathcal{N}\!\Big(b+\sum_{j=1}^{m}W^{(h_{j})}_{:,j},\,I\Big),$
so each chosen state contributes a single template vector and the mean is a sum of selected templates. This preserves the RBM’s locally linear structure encoding the continuous variables using the set of latent discrete variables. A key advantage of the Potts node is that each slot has its own weight vector, so swapping slots does not alter the previous slot’s learned weight. In practice, this yields sharper posteriors and more degrees of freedom than simply enforcing a one-hot encoding across multiple Bernoullis, which intrinsically lack intra-layer coupling.
GM–RBM is a modification where we replace binary hidden units with categorical slots but retain Gaussian visibles and standard RBM conditionals. We assess GM–RBM on hetero-associative recall and image modeling. To highlight the intrinsic strength of the GMRBM to the GBRBM, we develop comparison protocols that are capacity-matched (hidden nodes) and parameter-matched (weights), which appear later.
3 Theoretical Foundations of the Gaussian–Multinoulli RBM (GM-RBM)
The Gaussian–Multinoulli RBM (GM-RBM) replaces binary hidden units with discrete $q$-state categorical variables while keeping a continuous visible layer. This yields a combinatorial latent space with simple, closed-form conditionals.
3.1 Notation
Let $v\in\mathbb{R}^{n}$ be the visible vector. The hidden code is $h=(h_{1},\dots,h_{m})$ with $h_{j}\in\{1,\dots,q\}$. Parameters are: visible bias $b\in\mathbb{R}^{n}$; hidden bias $c_{j,k}\in\mathbb{R}^{m}$; and state-specific templates $W^{(k)}_{:,j}\in\mathbb{R}^{n}$. Define the conditional mean
$\mu(h)=b+\sum_{j=1}^{m}W^{(h_{j})}_{:,j}$
3.2 Energy, Joint, and Conditionals
The energy is
$E(v,h)=\tfrac{1}{2}\sum_{i=1}^{n}(v_{i}-b_{i})^{2}-\sum_{j=1}^{m}c_{j,h_{j}}-\sum_{i=1}^{n}\sum_{j=1}^{m}W^{(h_{j})}_{ij}\,v_{i}.$
Completing the square gives $E(v,h)=\tfrac{1}{2}\lVert v-\mu(h)\rVert_{2}^{2}+K(h)$ with $K(h)=\tfrac{1}{2}(\lVert b\rVert_{2}^{2}-\lVert\mu(h)\rVert_{2}^{2})-\sum_{j}c_{j,h_{j}}$. The joint is $p(v,h)\propto\exp(-E(v,h))$. The conditionals are
$p(v\mid h)=\mathcal{N}\big(\mu(h),\bm{1}\big)$
$\qquad p(h_{j}=k\mid v)=\frac{\exp\big(c_{j,k}+(W^{(k)}_{:,j})^{\top}v\big)}{\sum_{k^{\prime}=1}^{q}\exp\big(c_{j,k^{\prime}}+(W^{(k^{\prime})}_{:,j})^{\top}v\big)}=\text{Softmax}{(c_{j,k^{\prime}}+(W^{(k^{\prime})}_{:,j})^{\top}v\big)}$
3.3 Architecture and Special Cases
Each slot contributes one of $q$ templates, so the codebook $\{\mu(h):h\in\{1,\dots,q\}^{m}\}$ has size $q^{m}$. When $q=2$ and parameters are tied as $W^{(1)}_{:,j}-W^{(2)}_{:,j}=\widetilde{W}_{:,j}$ and $c_{j,1}-c_{j,2}=\widetilde{b}_{j}$ with a corresponding recentering of $b$, the GM-RBM reduces to a Gaussian–Bernoulli RBM. There is no requirement that $q$ be even; experiments may vary $q$ according to task. It is also possible to use usual variance reweighting as opposed to unit variance used in the GB-RBM.
3.4 Parameter Count and Fair Comparison Protocols
GM-RBM has $n$ parameters for $b$, plus $mq$ for $c$, plus nmqnmq for $W$, for a total of n+mq(1+n)n+mq\,(1+n). A GB-RBM with $m^{\prime}$ binaries has $n+m^{\prime}+nm^{\prime}$. We evaluate under two protocols:
  • Parameter-matched: Match the number of latent assignments by setting $m^{\prime}\approx m\,\log_{2}q$, since $|\mathcal{H}_{\text{GM}}|=q^{m}$ and $|\mathcal{H}_{\text{GB}}|=2^{m^{\prime}}$. Essentially, choose $m^{\prime}$ so total parameters are comparable across models
  • Capacity-matched: Choose $|\mathcal{H}_{\text{GB}}|$ in such a way that it matches the total number of learned representation for a given $|\mathcal{H}_{\text{GM}}|$
These regimes separate architectural effects (slot exclusivity and templating) from raw codebook size.
3.5 Learning and Negative-phase Sampling
The log-likelihood gradient satisfies
$\tfrac{\partial}{\partial\theta}\log p(v)=\mathbb{E}_{p(h\mid v)}\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big]-\mathbb{E}_{p(v\mid h)}\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big],\quad\theta\in\{b,c,W\}.$
We approximate the model expectation with short Markov chains.
Block Gibbs.
Alternate $h\sim p(h\mid v)$ using the per-slot softmax and $v\sim p(v\mid h)=\mathcal{N}(\mu(h),\mathbf{1})$. The visible draw is exact and parameter free.
Gibbs with visible Langevin.
Some implementations replace the exact Gaussian draw with an unadjusted Langevin step using $\nabla_{v}\log p(v\mid h)=\mu(h)-v$ (Liao et al., 2022):
$v_{t+1}=v_{t}+\tfrac{\varepsilon^{2}}{2}\big(\mu(h_{t})-v_{t}\big)+\varepsilon\,\xi_{t},\quad\xi_{t}\sim\mathcal{N}(0,\mathbf{1}),\ \varepsilon>0.$
This introduces a stepsize–dependent discretization error; for fixed $h_{t}$, multiple Langevin steps only approximate the exact conditional. In many Gaussian RBM variants, each visible dimension is allowed its own variance, so that $p(v\mid h)=\mathcal{N}(\mu(h),\Sigma)$ with a diagonal covariance $\Sigma=\operatorname{diag}(\sigma^{2})$. In this case, the gradient becomes $\Sigma^{-1}(\mu(h_{t})-v_{t})$, and the drift term is modified to $(\mu(h_{t})-v_{t})/\sigma^{2}$ (applied elementwise). Intuitively, inserting a visible Langevin move between hidden-layer updates allows hidden units to exchange information indirectly through the visible layer, encouraging more cooperative encodings and improving the model’s ability to capture complex dependencies.
In contrast to the expensive Langevin update, we use exact block Gibbs updates for the GM-RBM. The reason we utilize this less expensive block Gibbs update is that the visible Langevin step is merely an approximate sampler for the same conditional distribution $p(v\mid h)$. It does not inject additional information into the model, but instead introduces stepsize-dependent bias at an extra computational cost. Meanwhile, the multi-state multinoulli latent variables already enable this information to be represented and shared across hidden units through standard block Gibbs updates, making an additional visible Langevin refinement unnecessary in our setting. To ensure fair comparison to GB-RBM baselines, we explicitly report whether a visible Langevin step is used, the stepsize, the number of steps, and whether chains are persistent.
Sampler cost, mixing, and our choice.
The visible Langevin variant adds at least one extra update and a stepsize hyperparameter per negative step, which increases compute relative to an exact Gaussian draw. In our setting the Gibbs visible update samples exactly from $N(\mu(h),\mathbf{1})$ with a single noise vector and no stepsize tuning. The GM-RBM did was able to achieve quality results without resorting to the more expensive Langevin update. This choice also highlights a modeling claim of this work: with slot-wise categorical latents, GM-RBM exhibits fast mixing under basic Gibbs updates without relying on heavier samplers. We report chains and effective sample diagnostics to substantiate this point.
3.6 Key Properties
  • Locally linear: given $h$, $p(v\mid h)$ is Gaussian with fixed covariance.
  • Globally discrete: the means form a finite codebook indexed by discrete slots.
  • Modular: slots contribute additively in $\mu(h)$ and independently in $p(h\mid v)$.